#29 Is there a relation between murder and arson? A random sample of 15 Midwest cities...

#29

Is there a relation between murder and arson? A random sample of 15 Midwest cities (over 10,000 population) gave the following information about annual number of murder and arson cases. (Reference: Federal Bureau of Investigation, U.S. Department of Justice.)

City 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
Murder 11 18 14 15 13 10 9 8 12 6 17 22 20 21 24
Arson 37 73 79 18 29 43 31 24 12 14 133 139 47 98 138

Use a 1% level of significance to test the claim that there is a monotone-increasing relationship between the ranks of murder and arson.

(a) Rank-order murder using 1 as the largest data value. Also rank-order arson using 1 as the largest data value. Then construct a table of ranks to be used for a Spearman rank correlation test.

City Murder
Rank x
Arson
Rank y
d = x - y d2
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15




























































Σd2 =

(c) Compute the sample test statistic. (Use 3 decimal places.)
Answer #1

a)

x-y=
x rank y rank d d2
11 9 2 4
5 6 -1 1
8 5 3 9
7 13 -6 36
9 11 -2 4
12 8 4 16
13 10 3 9
14 12 2 4
10 15 -5 25
15 14 1 1
6 3 3 9
2 1 1 1
4 7 -3 9
3 4 -1 1
1 2 -1 1
∑d2 130

c)

test statistic = rs=1-6Σd2/(n(n2-1))= 0.768